# Efficient Redundancy Reduction for Open-Vocabulary Semantic Segmentation

### v1

@@ III-C3 Impact on latency
III-D Efficient Semantic Context Fusion
It is well known that context is essential for semantic understanding [5, 32, 8]. Hence, ERR-Seg incorporates the Efficient Semantic Context Fusion (ESCF) to integrate spatial-level and class-level contextual information via the Spatial-level Fusion Module (SFM) and Class-level Fusion Module (CFM), respectively. Additionally, Spatial-level Sequence Reduction (SSR) and Class-level Sequence Reduction (CSR) strategies are introduced to accelerate the context fusion process without compromising performance.
III-D1 Spatial-level Fusion Module
Spatial-level Fusion Model (SFM) aims to model the relationship between different spatial locations within the same class. The architecture of SFM is presented in Fig. 5(a). Inspired by [52], SFM introduces Spatial-level Sequence Reduction (SSR) to reduce the computational complexity of the attention mechanism. Specifically, it reduces the sequence length of $K$ and $V$ by a factor of $R$, as follows:
$Z^{\prime}$ $=\mathrm{Reshape}\left(Z,\left(N/R,C\times R\right)\right),$ (20)
$Z$ $=\mathrm{Linear}\left(Z^{\prime},\left(C\times R,R\right)\right),$ (21)
where $Z\in\{K,V\}$, $N$ is the sequence length, $\mathrm{Reshape(.)}$ refers to reshaping $Z$ into the shape of $\left(N/R,C\times R\right)$, and $\mathrm{Linear(.)}$ represents a fully connected layer that projects the input from $C\times R$ dimension to $C$ dimension. As a result, it merges adjacent $R$ embeddings into one embedding and reduces the complexity of the self-attention mechanism from $O(N^{2})$ to $O(N^{2}/R)$. The cost maps $S_{r}$ will then go through a depth-wise convolution layer, followed by a two-layer MLP. Moreover, we replace the first linear layer of the MLP with a star operation [53] to leverage its ability to map to high-dimensional spaces. More specifically, the star operation is defined as follows:
$S_{r}^{(1)}$ $=\mathrm{Linear}\left(S_{r},\left(C,C^{\prime}\right)\right),$ (22)
$S_{r}^{(2)}$ $=\mathrm{Linear}\left(S_{r},\left(C,C^{\prime}\right)\right),$ (23)
$S_{r}$ $=\sigma\left(S_{r}^{(1)}\right)*S_{r}^{(2)},$ (24)
where $\sigma$ denotes the activation function, and $C^{\prime}$ is the intermediate embedding dimension of the MLP.
III-D2 Class-level Fusion Module
Class-level fusion module (CFM) aims to capture dependencies within the same spatial location but across different classes. The architecture of our proposed CFM is depicted in Fig. 5(b). Noticing that surrounding pixels tend to exhibit similar features, CFM introduces the Class-level Sequence Reduction (CSR), which employs the average pooling to reduce the size of the cost maps and residual connections to incorporate the upsampling results and origin input.
III-E Gradual Upsampling Decoder
After the ESCF, low-resolution semantic maps $E_{r}\in\mathbb{R}^{P\times C\times H_{c}\times W_{c}}$ are obtained. To recover high-resolution semantic maps, we propose the Gradual Upsampling Decoder (GUD), leveraging the high-rank details from $\mathcal{F}_{i}^{v},i\in\{1,2,3\}$ to guide the upsampling process of $E_{r}$. The architecture of GUD is illustrated in Fig. 3. Initially, $\mathcal{F}_{i}^{v},i\in\{1,2,3\}$ undergoes a $3\times 3$ convolution and is then concatenated with the cost maps at each class. The concatenated features are then passed through the guidance refine block, comprising two $3\times 3$ convolutions. Our observations indicate that a small kernel yields better performance, as the texture is local and does not require a large receptive field. Moreover, we emphasize the significance of stopping the gradient of $\mathcal{F}_{i}^{v},i\in\{1,2,3\}$ as it can prevent the visual encoder from distribution shift.
IV Experiments
TABLE IV: Label-level similarity between evaluation and training datasets.
Dataset $1.00$ $\geq 0.95$ $\geq 0.90$
PAS-20 0.45 0.90 1.00
PC-59 0.30 0.85 0.95
A-150 0.19 0.41 0.57
PC-459 0.10 0.21 0.29
A-847 0.06 0.12 0.21
TABLE V: Component Analysis. We remove modules from ERR-Seg independently to evaluate their impact.
Module A-847 PC-459 A-150 PC-59 PAS-20
ERR-Seg 13.9 20.5 35.8 56.9 95.4
w/o HSM 13.3 19.3 35.5 56.4 95.0
w/o CRM 14.0 20.1 35.3 56.8 94.9
w/o ESCF 9.3 12.8 26.1 43.6 90.8
w/o GUD 13.4 19.7 34.8 56.4 95.1
TABLE VI: Ablation study for HSM. Different gradient sources of updating stage 4 of CLIP’s visual encoder are tried.
@@ IV Experiments
IV-A Experimental Setup
IV-A1 Traning dataset
Following the existing OVSS methods [16, 15], we train our model on COCO-Stuff [54]. It has 118,287 images for training with 171 different classes.
IV-A2 Evaluation dataset
  • PASCAL-Context [55] contains 459 classes with 5,104 images for validation. We report the scores of the full classes (PC-459) and the scores of the 59 most frequent classes (PC-59).
  • ADE20K [56] has 2,000 validation images with 847 classes. Similarly to PASCAL-Context, we report the results of 150 frequent classes (A-150) and full 847 classes (A-847).
  • PASCAL VOC [57] includes 1,449 images for validation with 20 common object classes. We report the performance of the 20 classes named PAS-20.
IV-A3 Analysis of the evaluation dataset
Prior studies have typically overlooked the label-level similarity between the validation and training sets, while it is necessary for open vocabulary-related tasks. In this section, we extract text embeddings for each category and then calculate the cosine similarity between the categories in the evaluation and training datasets, resulting in a similarity matrix. Subsequently, we map each category in the evaluation dataset to a category with the highest similarity score in the training dataset. The results are presented in Table IV; we report the proportion of labels in the validation sets with a similarity score equal to $1.00$, greater than $0.95$, $0.90$, and $0.85$.
IV-A4 Implementation details
We adopt the Convolutional CLIP [58] from OpenCLIP [59], where the visual encoder is ConvNeXt [60] architecture. Our code is implemented on the top of Detectron2 [61]. The AdamW [62] optimizer is employed, with an initial learning rate of $2\times 10^{-4}$ and a weight decay of $1\times 10^{-4}$. The learning rate is set to $2\times 10^{-6}$ for fine-tuning the CLIP model. Our model is trained on COCO-Stuff using cross-entropy loss for $80$K iterations, with a batch size of $4$ and a crop size of $640\times 640$. $4$ NVIDIA RTX 3090 GPUs were used for training.
Figure 6: Impact of score coefficient $\lambda$ on CRM.
IV-B Comparing with State-of-the-art
We compare our proposed ERR-Seg with existing state-of-the-art methods on the five standard evaluation settings, with the results presented in Table II. To ensure a fair comparison, all models utilize CLIP as the VLM. We also include whether CLIP’s image and text encoders are fine-tuned. When both the image and text encoders are fine-tuned, ERR-Seg achieves remarkable performance on these five settings, particularly in the base model. Notably, on A-847, PC-459, A-150 and PAS-20, ERR-Seg exhibited significant improvements, outpacing CAT-Seg [16] by $15.8\%$, $7.9\%$, $12.6\%$ and $0.8\%$, respectively. In the large model, ERR-Seg surpasses the previous arts on A-847, PC-459 and A-150 by $5.6\%$, $0.4\%$ and $1.1\%$, respectively. When only the text encoder is fine-tuned, ERR-Seg also outperforms SED [15] on all five settings, with an improvement of $13.7\%$, $0.4\%$, $5.2\%$, $0.2\%$ and $0.6\%$ on A-847, PC-459, A-150, PC-59 and PAS-20, respectively.
However, ERR-Seg does not perform as well as CAT-Seg on PC-59, exhibiting a $4.1\%$ lower performance than CAT-Seg in the large model. As discussed in Section IV-A3, PC-59 primarily assesses in-domain open-vocabulary performance. This suggests that CAT-Seg might be overfitting to the seen categories during training, thus showcasing superior in-domain performance compared to our model but lagging in out-domain performance. In contrast, with the introduction of CRM, ERR-Seg customizes the vocabulary list for each image during training, which helps prevent overfitting to the seen categories.
Table III presents the results of the latency comparisons. Although cost-based CAT-Seg shows notable performance enhancement over mask-based SAN, it also results in a considerable increase in latency. In contrast, our proposed ERR-Seg not only further improves the performance of cost-based methods, but also reduces its latency significantly. Specifically, the latency is reduced by $79.3\%$ on A-847 and $77.0\%$ on A-150.
TABLE VII: Effect of Top $k$ categories on CRM.
@@ IV-B Comparing with State-of-the-art
TABLE VIII: The impact of Spatial-level Fusion Module (SFM) and Class-level Fusion Module (CFM).
ESCF A-847 PC-459 A-150 PC-59 PAS-20
SFM 13.1 19.2 34.4 56.3 94.4
CFM 13.0 19.6 34.1 55.4 95.5
Both 13.9 20.5 35.8 56.9 95.4
TABLE IX: Impact of star operation.
ESCF A-847 PC-459 A-150 PC-59 PAS-20
w/o star 14.1 20.1 35.6 56.3 94.7
with star 13.9 20.5 35.8 56.9 95.4
TABLE X: Effect of the Spatial-level Sequence Reduction (SSR) and Class-level Sequence Reduction (CSR).
SSR CSR A-847 PC-459 A-150 PC-59 GFLOPs Memory
13.5 20.0 35.4 57.0 188.1 9.5
14.2 20.0 35.9 56.9 148.9 7.8
14.2 19.9 35.6 57.0 134.6 5.9
13.9 20.5 35.8 56.9 95.3 4.2
IV-C Ablation Study
IV-C1 Component analysis
The ERR-Seg framework comprises four main branches: Hierarchical Semantic Module (HSM), Channel Reduction Module (CRM), Efficient Semantic Context Fusion (ESCF), and Gradual Upsampling Decoder (GUD). To assess the importance of these modules, each module is individually removed, and the results are shown in Table V. The findings demonstrate the contribution of ESCF and GUD in enhancing performance, with improvements of +$9.7$ and +$1.0$ mIoU on A-150, respectively. A more detailed efficiency analysis of the ESCF will be provided later. Without using CRM, the redundancy reduction process is eliminated during training and evaluation, resulting in a decrease of $0.4$ mIoU on PC-459 and $0.5$ mIoU on A-150. Additionally, HSM yields a +$1.2$ mIoU enhancement on PC-459 and a +$0.5$ mIoU improvement on A-150, indicating the significance of hierarchical cost maps.
IV-C2 Ablation study for HSM
Without HSM, the gradient for updating stage 4 of CLIP’s visual encoder is derived solely from $\mathcal{F}_{4}^{s}$. However, when HSM is active, the gradient can be sourced from $\mathcal{F}_{3}^{s}$ and $\mathcal{F}_{4}^{s}$, as HSM functions as a parallel branch. Table VI investigates the impact of different gradient sources on updating stage 4. The results clearly show that the combination of gradients from $\mathcal{F}_{3}^{s}$ and $\mathcal{F}_{4}^{s}$ yields superior outcomes compared to utilizing either source individually. Furthermore, relying solely on $\mathcal{F}_{3}^{s}$ leads to notably poorer model performance, since $\mathcal{F}_{4}^{s}$ represents the original CLIP output and offers more precise semantic details.
Figure 8: Visualization of segmentation results on ADE20K-150 and ADE20K-847. The results of previous state-of-the-art methods SAN [15] and CAT-Seg [16] are also included for comparison.
Figure 9: Visualization of segmentation results in various domains. Our proposed ERR-Seg is capable of segmenting capybaras (a rare category in public datasets) from various domains, including (a) synthesized images, (b) cartoon images, (c) natural images, and (d) capybara dolls. Moreover, ERR-Seg achieves more precise masks than SAN [14] and CAT-Seg [16].
IV-C3 Impact of different design of CRM
This section investigated the impact of hyperparameters within CRM. Fig. 6 illustrates how the score coefficient $\lambda$ influences class scoring, with optimal performance achieved at $\lambda=0.1$. A larger $\lambda$ may blur the distinction between classes with similar rankings, while a smaller $\lambda$ might overlook certain classes. Furthermore, Fig. 7 explores the effect of the selected class number $P$ during training on the COCO-Stuff dataset, highlighting that setting $P=32$ results in the highest performance. It also reveals that setting $P$ either too large or too small can hinder the modeling of contextual information. Table VII presents the effect of the top $k$ categories of a pixel, indicating that the CRM is minimally sensitive to $k$, with slightly superior results observed at $k=3$. A smaller $k$, such as 1, risks overlooking existing categories due to potential inaccuracies in cost maps’ prior knowledge. Conversely, a larger $k$ may introduce redundant categories, hindering CRM’s ability to identify the most relevant classes.
IV-C4 Ablation study for ESCF
ESCF aims to incorporate more contextual information at both the class and spatial levels within the cost maps. We observed a significant decrease in performance when the Spatial-level Fusion Module (SFM) and the Class-level Fusion Module (CFM) were removed, as shown in Table VIII. Additionally, Table IX presents the results obtained without the star operation but MLP instead. It shows that the performance of ERR-Seg is improved on four datasets through the star operation. Table X presents the effects of the sequence reduction strategies, SSR and CSR, including their impact on GFLOPs and memory usage (GiB). The findings indicate that the use of CSR or SSR alone does not negatively affect performance. Overall, ESCF achieves a reduction of $49.6\%$ GFLOPs and $55.6\%$ memory without compromising performance.
TABLE XI: Ablation study for GUD
Guidance Feature A-847 PC-459 A-150 PC-59 PAS-20
$\mathcal{F}_{1}^{v},\mathcal{F}_{2}^{v}$ 14.0 20.3 35.3 57.1 95.3
$\mathcal{F}_{1}^{v},\mathcal{F}_{2}^{v},\mathcal{F}_{3}^{v}$ 13.9 20.5 35.8 56.9 95.4
Guidance Gradient A-847 PC-459 A-150 PC-59 PAS-20
w/o stop 13.9 19.7 35.1 56.9 94.9
with stop 13.9 20.5 35.8 56.9 95.4
Kernel Size A-847 PC-459 A-150 PC-59 PAS-20
3 13.9 20.5 35.8 56.9 95.4
5 13.8 20.3 35.4 56.6 95.2
7 13.7 20.1 35.7 56.7 95.1
IV-C5 Ablation study for GUD
GUD utilizes the middle-layer features ($\mathcal{F}_{i}^{v},1\leq i\leq 3$) from the CLIP’s visual encoder to reconstruct high-resolution segmentation results. We first investigate the necessity of integrating $\mathcal{F}_{3}^{v}$, as it shares the same size as the original cost maps. The experimental results are shown in the top section of Table XI, demonstrating the effectiveness of incorporating with $\mathcal{F}_{3}^{v}$, resulting in a +$0.5$ mIoU improvement on A-150. Moreover, we investigate the impact of updating the visual encoder with the gradient of $\mathcal{F}_{i}^{v}(1\leq i\leq 3)$, as this could potentially disrupt the alignment between the visual and text encoders. The results are presented in the middle section of Table XI. It shows that preventing the gradient of $\mathcal{F}_{i}^{v}(1\leq i\leq 3)$ from updating the visual encoder leads to better performance. The lower section of Table XI presents the effect of convolution kernel size, revealing that a smaller kernel size improves performance due to the localized nature of textual information.
TABLE XII: Impact of CLIP Architecture
CLIP Arch. A-847 PC-459 A-150 PC-59 PAS-20
ViT-B/16 7.5 13.0 22.8 49.7 91.0
ConvNeXt-B 9.7 12.5 26.9 49.4 89.9
TABLE XIII: The impact of sampled class number of CRM during inference.
@@ IV-C5 Ablation study for GUD
IV-C6 Impact of CLIP architecture

### v2

@@ 4.1.4 Implementation details
4.2 Main Results
4.2.1 Comparing with state-of-the-art
We compare our proposed ERR-Seg with existing state-of-the-art methods on the five standard evaluation settings, with the results presented in Table 3. To ensure a fair comparison, all models utilize CLIP as the VLM. We also include whether CLIP’s image and text encoders are fine-tuned. When both the image and text encoders are fine-tuned, ERR-Seg achieves remarkable performance on these five settings, particularly in the base model. Notably, on A-847, PC-459, and A-150, ERR-Seg achieves significant gains, surpassing the previous best results by margins of $5.3\%$, $7.9\%$, and $11.9\%$, respectively. In the large model setting, ERR-Seg continues to outperform prior arts, exceeding the best previous scores on A-847, PC-459, and A-150 by $4.3\%$, $0.4\%$, and $1.1\%$, respectively. When only the text encoder is fine-tuned, ERR-Seg also outperforms SED [xie2024sed] on all five settings, with an improvement of $13.7\%$, $0.4\%$, $5.2\%$, $0.2\%$ and $0.6\%$ on A-847, PC-459, A-150, PC-59 and PAS-20, respectively.
However, ERR-Seg does not perform as well as CAT-Seg on PC-59, exhibiting a $4.1\%$ lower performance than CAT-Seg in the large model. As discussed in Section 4.1.3, PC-59 primarily assesses in-domain open-vocabulary performance. This indicates that CAT-Seg might fit more of the seen categories during training, thus showcasing superior in-domain performance compared to our model but lagging in out-domain performance. In contrast, with the introduction of redundant class reduction, ERR-Seg customizes the vocabulary list for each image during training, which helps prevent overfitting to the seen categories. A similar trend is observed with SCAN: it outperforms CAT-Seg by 0.9 and 2.4 on PC-59 and PAS-20, respectively, but suffers a significant performance drop of 8.2 on PC-459.
Table 5: Comparison on seen and unseen classes.
Method A-847 PC-459 A-150
Seen Unseen Seen Unseen Seen Unseen
CAT-Seg 29.2 9.6 38.3 12.5 42.1 24.4
ERR-Seg 30.1 11.6 40.1 13.9 44.7 29.4
Tab. 5 presents a comparison with CAT-Seg on both seen and unseen classes. Here, unseen classes are defined as those with a label similarity exceeding 0.95 to the COCO-Stuff training set, based on the analysis in Sec. 4.1.3. The results demonstrate that our method not only generalizes effectively from seen to unseen classes, but also achieves consistent improvements across both settings. Furthermore, our method exhibits more substantial gains on unseen classes relative to CAT-Seg: on the seen classes of A-847, PC-459, and A-150, the relative improvements are $3.1\%$, $4.7\%$, and $6.2\%$, respectively, while on unseen classes, the relative improvements reach $20.8\%$, $11.2\%$, and $20.5\%$.
4.2.2 Efficiency comparison
Table 4 presents the results of the latency comparisons. Although cost-based CAT-Seg shows notable performance enhancement over mask-based SAN, it also results in a considerable increase in latency. In contrast, our proposed ERR-Seg not only further improves the performance of cost-based methods but also reduces their latency significantly. Specifically, the latency is reduced by $79.3\%$ on A-847 and $77.0\%$ on A-150.
4.2.3 Qualitative results
Fig. 5 presents a qualitative comparison among SAN [xu2023side], CAT-Seg [cho2024cat], and our proposed ERR-Seg on the validation sets of ADE20K-150 and ADE20K-847. The results highlight the superiority of our method over previous approaches in mask generation and open-vocabulary recognition. In addition, ERR-Seg can identify targets that were previously overlooked in ADE20K. While the current OVSS model shows remarkable open-vocabulary recognition capability due to the advancements of VLMs, it struggles with segmenting intricate masks. There is still potential for improvement in effectively recognizing small and nested targets.
4.3 Ablation Study and Analysis
4.3.1 Ablation study on RRHC
Table 6 presents the ablation study on RRHC. The results indicate that removing the class redundancy leads to performance increases on some benchmarks, such as +0.4+0.4 mIoU on PC-459 and +0.5+0.5 mIoU on A-150. This demonstrates that our proposed class redundancy removal strategy can serve not only as a methodology to alleviate computation overload for OVSS but also to improve performance. Moreover, it also demonstrates the effectiveness of our hierarchical cost maps, which leads to +1.2+1.2 mIoU improvement on PC-459.
Without RRHC, the gradient for updating stage 4 of CLIP’s visual encoder is derived solely from $\mathcal{F}_{4}^{s}$. However, when using RRHC, the gradient can be sourced from $\mathcal{F}_{3}^{s}$ and $\mathcal{F}_{4}^{s}$, as it functions as a parallel branch. Table 6 investigates the impact of different gradient sources on updating stage 4. The results show that relying solely on $\mathcal{F}_{3}^{s}$ leads to notably poorer model performance. Furthermore, the combination of gradients from $\mathcal{F}_{3}^{s}$ and $\mathcal{F}_{4}^{s}$ yields slightly better performance on some benchmarks than relying solely on $\mathcal{F}_{4}^{s}$.
Table 6: Ablation study and on RRHC.
Class redundancy A-847 PC-459 A-150 PC-59 PAS-20
$w/o$ remove 14.0 20.1 35.3 56.8 94.9
$w/$ remove 13.9 20.5 35.8 56.9 95.4
Cost maps A-847 PC-459 A-150 PC-59 PAS-20
Single 13.3 19.3 35.5 56.4 95.0
Hierarchical 13.9 20.5 35.8 56.9 95.4
@@ 4.3.1 Ablation study on RRHC
4.3.2 Hyperparameter analysis for RRHC
Fig. 6 illustrates how the score coefficient $\lambda$ influences class scoring, with optimal performance achieved at $\lambda=0.1$. A larger $\lambda$ may blur the distinction between classes with similar rankings, while a smaller $\lambda$ might overlook certain classes.
Table 7 presents the effect of the top $k$ categories of a pixel, indicating that the redundant class reduction mechanism is minimally sensitive to $k$, with slightly superior results observed at $k=3$. A smaller $k$, such as 1, risks overlooking existing categories due to potential inaccuracies in cost maps’ prior knowledge. Conversely, a larger $k$ may introduce redundant categories, hindering its ability to identify the most relevant classes.
Fig. 7 explores the effect of the selected class number $P$ during training on the COCO-Stuff dataset, highlighting that setting $P=32$ yields optimal performance. It also reveals that setting $P$ either too large or too small can hinder the modeling of contextual information. During inference, $P$ is set to the original number of classes of each dataset to ensure a fair comparison.
Table 8 illustrates the influence of the sampled class number $P$ in the redundant class reduction mechanism during the inference phase. It is evident that as $P$ increases gradually, the performance tends to reach a saturation point. The performance is observed to saturate as $P$ increases, consistent with the expectation that a sufficiently large $P$ can encompass all relevant categories. Specifically, we assign the following values to $P$: 48 for A-847, 48 for PC-459, 32 for A-150, 24 for PC-59, and 16 for PAS-20.
Table 7: Effect of Top $k$ categories.
@@ 4.3.2 Hyperparameter analysis for RRHC
Figure 6: Impact of score coefficient $\lambda$.
Table 8: The impact of sampled class number $P$ during inference.
@@
4.3.3 Analysis on RRHC
Table 9 presents the impact of removing redundant classes in RRHC on the final segmentation performance. We report the average number of classes per image in the ground truth, as well as the number of predicted classes overlapping with the ground truth, both before and after redundant class removal. It can be observed that the removal leads to a slight reduction in the number of recognized classes, for instance, on A-150, an average of 0.2 classes per image is no longer detected. Nevertheless, this minor decline is a worthwhile trade-off, as removing redundant classes brings gains in both segmentation accuracy (e.g., 0.4 and 0.5 improvement on PC-459 and A-150, respectively) and inference speed (discussed in Sec. 3.4).
Table 9: Analysis of redundant class removal in RRHC. The “#Classes” columns denote the average number of predicted classes that overlap with the ground truth per image.
Method A-847 PC-459 A-150
#Classes mIoU #Classes mIoU #Classes mIoU
Ground Truth 10.2 - 6.5 - 8.5 -
$w/o$ removal 6.2 14.0 4.7 20.1 6.1 35.3
$w/$ removal 5.7 13.9 4.4 20.5 5.9 35.8
4.3.4 Ablation study on RRCA
RRCA aims to incorporate more contextual information at both the class-level and spatial-level within the cost maps. We first conduct experiments to analyze the spatial-level and class-level aggregation separately. We observed a significant decrease in performance when the spatial-level aggregation and the class-level aggregation are removed, as shown in Table 10. Table 11 presents the effects of the sequence reduction strategies, SSR and CSR, including their impact on GFLOPs, memory usage (GiB) and latency (ms). The findings indicate that the use of CSR or SSR alone does not negatively affect performance. Overall, equipping SSR and CSR achieves a reduction of $49.6\%$ GFLOPs, $55.6\%$ memory and $52.4\%$ latency without compromising performance. Additionally, Table 12 compares the effect of different MLP architectures within the cost aggregation blocks. Here, Vanilla refers to the standard MLP consisting of two linear layers and an activation function, while the Modified MLP incorporates the star operation to strengthen its representational capacity. It shows that the performance of ERR-Seg is improved on four datasets by adopting the modified MLP.
Table 10: The impact of spatial-level aggregation and class-level aggregation.
Spatial Class A-847 PC-459 A-150 PC-59 PAS-20
9.3 12.8 26.1 43.6 90.8
13.1 19.2 34.4 56.3 94.4
13.0 19.6 34.1 55.4 95.5
13.9 20.5 35.8 56.9 95.4
Table 11: Effect of the Spatial-level Sequence Reduction (SSR) and Class-level Sequence Reduction (CSR).
SSR CSR A-847 PC-459 A-150 PC-59 GFLOPs Memory Latency
13.5 20.0 35.4 57.0 188.1 9.5 103
14.2 20.0 35.9 56.9 148.9 7.8 91
14.2 19.9 35.6 57.0 134.6 5.9 57
13.9 20.5 35.8 56.9 95.3 4.2 49
Table 12: Impact of different MLP structures in cost aggregation blocks.
MLP A-847 PC-459 A-150 PC-59 PAS-20
Vanilla 14.1 20.1 35.6 56.3 94.7
Modified 13.9 20.5 35.8 56.9 95.4
4.3.5 Impact of CLIP architecture